# bibliography.bib

### v1

@@ 4 Extensions of frame maps
Any frame morphism $f:L\rightarrow M$ between spatial frames lifts to a coframe morphism
$f_{\mathcal{CP}}:\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op}\rightarrow\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(M))^{op}.$
@@
Any frame morphism $f:L\rightarrow M$ between pre-spatial frames lifts to a coframe morphism
$f_{\mathcal{SO}}:\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op}\rightarrow\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(M))^{op}.$
@@
There is a forgetful functor $\pi_{1}:\mathbf{Raney}\rightarrow\mathbf{Frm}$ which forgets about the second component of the extension. We observe that the assignment $L\mapsto(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$ on objects can be extended to a functor $\mathsf{Filt}_{\mathcal{SE}}$ by mapping each frame morphism to the morphism $f_{\mathcal{SE}}$ whose existence is established by Lemma 4.4.
Theorem 4.8.
For a frame $L$, the pair $(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$ is the free Raney extension over it, that is, we have $\mathsf{Filt}_{\mathcal{SE}}\dashv\pi_{1}$. In particular, the category of frames is a full coreflective subcategory of $\mathbf{Raney}$.
Proof.
Suppose that we have a frame map $f:L\rightarrow M$. Let $(M,D)$ be a Raney extension. By Lemma 3.5, we have $D^{*}\subseteq\mathsf{Filt}_{\mathcal{SE}}(M)^{op}$, and by Lemma 4.4, preimages of strongly exact filters are strongly exact. Therefore, preimages of elements in $D^{*}$ are in $\mathsf{Filt}_{\mathcal{SE}}(L)$. By Theorem 4.3, the frame map lifts as required, and so we have $\mathsf{Filt}_{\mathcal{SE}}\dashv\pi_{1}$. ∎
We use the isomorphism in Theorem 2.6 to prove a universal property of the collection $\mathsf{S}_{\mathfrak{o}}(L)$ of fitted sublocales of a frame.
Corollary 4.9.
For a frame $L$, the pair $(L,\mathsf{S}_{\mathfrak{o}}(L))$ is the free Raney extension on it.
We already have seen that each frame has the largest Raney extension. Every frame also has the smallest Raney extension.
Lemma 4.10.
Proof.
Proposition 4.11.
For all Raney extensions $(L,C)$ we have a surjection
$(L,C)\rightarrow(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op}),$
$c\mapsto\mbox{cl}_{\mathcal{E}}({\uparrow}^{L}c).$
Proof.
The collection $C^{*}\subseteq\mathsf{Filt}(L)$ is a sublocale which contains all principal filters, and so $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}$ by Lemma 4.10. This can be rephrased as saying that the preimage map relative to the identity $1_{L}:L\to L$, which is itself the identity on $\mathsf{Filt}(L)$, maps exact filters to filters of $C^{*}$. Then by Theorem 4.3 we have a surjection $c\mapsto\mbox{cl}_{\mathcal{E}}({\uparrow}^{L}c)$. ∎
Corollary 4.12.
For a frame $L$, the smallest Raney extension on it is $(L,\mathsf{S}_{\mathfrak{c}}(L)^{op})$.
Theorem 4.13.
Proof.
That every Raney extension belongs to the section $[\mathsf{Filt}_{\mathcal{E}}(L),\mathsf{Filt}_{\mathcal{SE}}(L)]$ follows from Theorem 4.8 and Proposition 4.11. Suppose that we have a sublocale $\mathcal{F}\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)$ such that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathcal{F}$. By Lemma 4.10, $\mathcal{F}$ contains all principal filters, and so, by Theorem 3.6, the pair $(L,\mathcal{F}^{op})$ is a Raney extension. ∎
4.1 Functoriality of the exact filters extension
The assignment $L\mapsto(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ is not always functorial, as shown in [ball19], where the question is explored for structure $\mathsf{S}_{\mathfrak{c}}(L)$, isomorphic to $\mathsf{Filt}_{\mathcal{E}}(L)$. For a frame morphism $f:L\rightarrow M$, we will say that it is exact if whenever the meet of a family $\{x_{i}:i\in I\}\subseteq L$ is exact, so is the meet of $\{f(x_{i}):i\in I\}$, and furthermore $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})$.
Proposition 4.14.
A morphism $f:L\rightarrow M$ is exact if and only if preimages of exact filters are exact. This holds if and only if the morphism can be extended to a morphism
Proof.
Suppose that $f:L\rightarrow M$ is an exact frame map, and that $G\subseteq M$ is an exact filter. Suppose that $\bigwedge_{i}x_{i}\in L$ is an exact meet such that $f(x_{i})\in G$. By exactness of this map, the meet $\bigwedge_{i}f(x_{i})$ is exact and so $\bigwedge_{i}f(x_{i})\in G$. Again, by exactness of $f$, we have $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})$. Indeed, then, $\bigwedge_{i}x_{i}\in f^{-1}(G)$. Conversely, suppose that we have a frame map $f:L\rightarrow M$ such that it is not exact. This means that either we have an exact meet $\bigwedge_{i}x_{i}\in L$ such that it is not preserved by $f$, or we have an exact meet $\bigwedge_{i}x_{i}\in L$ such that $\bigwedge_{i}f(x_{i})$ is not exact. We consider these two cases in turn. In the first case, we consider the principal filter ${\uparrow}\bigwedge_{i}f(x_{i})$. This is exact, as it is closed under all meets. We notice that by our hypothesis $f(\bigwedge_{i}x_{i})$ is not an element of this filter. Let us call $F$ the preimage of this filter. We have that $x_{i}\in F$ but $\bigwedge_{i}x_{i}\notin F$, and so $F$ is not exact. In the second case, consider an exact meet $\bigwedge_{i}x_{i}\in L$ such that $\bigwedge_{i}f(x_{i})$ is not exact. In particular, let $y\in M$ be such that $\bigwedge_{i}(f(x_{i})\vee y)\nleq(\bigwedge_{i}f(x_{i}))\vee y$. We now consider the exact filter
That this is an exact filter follows from the characterization of Lemma 2.7. Let $F$ be the preimage of this filter. We have that $x_{i}\in F$ for each $i\in I$. We claim that $\bigwedge_{i}x_{i}\notin F$. This follows from the fact that by our hypothesis $\bigwedge_{i}(f(x_{i})\vee y)\nleq(\bigwedge_{i}f(x_{i}))\vee y$ and $f(\bigwedge_{i}x_{i})\leq\bigwedge_{i}f(x_{i})$. The rest of the claim follows by Theorem 4.3. ∎
Let us call $\mathbf{Frm}_{\mathcal{E}}$ the category of frames with exact maps. Let us also call $\mathbf{Raney}_{\mathcal{E}}$ the category of Raney extensions with morphisms that, restricted to the frame components, are exact. The assignment $L\mapsto(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$, then, determines a functor $\mathsf{Filt}_{\mathcal{E}}:\mathbf{Frm}_{\mathcal{E}}\to\mathbf{Raney}_{\mathcal{E}}$.
Theorem 4.15.
We have an adjunction $\pi_{1}:\mathbf{Raney}_{\mathcal{E}}\leftrightarrows\mathbf{Frm}_{\mathcal{E}}:\mathsf{Filt}_{\mathcal{E}}$ with $\pi_{1}\dashv\mathsf{Filt}_{\mathcal{E}}$.
Proof.
Suppose that we have an exact frame map $f:L\rightarrow M$, and that $(L,C)$ is a Raney extension. By Proposition 4.14, as $f$ is exact, preimages of filters in $\mathsf{Filt}_{\mathcal{E}}(M)$ are in $\mathsf{Filt}_{\mathcal{E}}(L)$. Furthermore, $C^{*}$ must contain all principal filters, and so by Lemma 4.10 this implies that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}$. Then, preimages of exact filters of $M$ are in $C^{*}$. By Theorem 4.3, then, we have a map of Raney extensions $(L,C)\rightarrow(M,\mathsf{Filt}_{\mathcal{E}}(M)^{op})$ extending $f$, as desired. ∎
5 Canonical extensions as Raney extensions
We now look at the notion of canonical extension of a frame from [jakl20], and see how it relates to Raney extensions. We will also look at how the Prime Ideal Theorem relates to some of our results. For a pre-spatial frame $L$, we will call its canonical extension the Raney extension
$(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op}).$
@@ 5 Canonical extensions as Raney extensions
This is indeed a Raney extension, by Proposition 2.10. For a Raney extension $(L,C)$, we say that an element $c\in C$ is compact if, for every directed collection $D\subseteq L$, we have that $c\leq\bigvee D$ implies that $c\leq d$ for some $d\in D$. We say that a Raney extension $(L,C)$ is algebraic if every element of $c$ is the join of compact elements.
@@
Notice that an element $x\in C$ is compact if and only if the filter ${\uparrow}^{L}x$ is Scott-open. Consider the isomorphism ${\uparrow}^{L}:C\cong C^{*}$. The Raney extension $(L,C)$ is algebraic if and only if in $C^{*}$ every element is a join of Scott-open filters of the form ${\uparrow}^{L}x$ for some $x\in L$. We have that $C^{*}\subseteq\mathsf{Filt}(L)^{op}$ is a subcolocale inclusion, and subcolocale inclusions preserves all joins, and joins in $\mathsf{Filt}(L)^{op}$ are intersections. Therefore, algebraicity of $(L,C)$ is equivalent to every filter in $C^{*}$ being an intersection of Scott-open filters in $C^{*}$. ∎
@@
For a pre-spatial frame $L$, its canonical extension is the free algebraic Raney extension over it. That is, whenever we have a frame map $f:L\rightarrow M$ and $(M,C)$ is an algebraic Raney extension, the map $f$ can be extended to a map of Raney extensions
$(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})\rightarrow(M,C).$
@@
The assignment is functorial, by Lemma 4.7. Suppose that $L$ is a pre-spatial frame, and that $(M,C)$ is an algebraic Raney extension. Suppose that we have a frame map $f:L\rightarrow M$. Consider the canonical extension $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})$. We have that, as $(M,C)$ is algebraic, $C^{*}\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(M))^{op}$, by Lemma 5.1. By Lemma 4.7, preimages of Scott-open filters are Scott-open. Then, preimages of filters in $C^{*}$ are in $\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$. By Theorem 4.3 this means that there is a map of Raney extensions $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})\rightarrow(M,C)$ extending the frame map $f:L\rightarrow M$. ∎
We now look at the role of the Prime Ideal Theorem, and work towards characterizing sobriety and strict sobriety of spaces.

### v2

@@ 6.3 Free spatial Raney extension on a spatial frame
Any frame morphism $f:L\rightarrow M$ between spatial frames extends to a Raney morphism
$f_{\mathcal{CP}}:(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})\rightarrow(M,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(M))^{op}).$
The lemma above shows that there is a functor $\mathsf{Filt}_{\mathcal{CP}}:\mathbf{spFrm}\to\mathbf{Raney}$ from the category of spatial frames, assigning to a frame $L$ the Raney extension $(L,\mathsf{Filt}_{\mathcal{CP}}(L)^{op})$, and to a morphism $f:L\to M$ the Raney map $f_{\mathcal{CP}}:(L,\mathsf{Filt}_{\mathcal{CP}}(L)^{op})\to(M,\mathsf{Filt}_{\mathcal{CP}}(M)^{op})$ extending it.
Proposition 6.5.
For a spatial frame $L$, the pair $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})$ is the free spatial Raney extension over it. In particular, the category of spatial frames is a full coreflective subcategory of that of spatial Raney extensions.
Proof.
Suppose that $f:L\rightarrow M$ is a map between spatial frames, and that $(M,C)$ is a spatial Raney extension. By spatiality, we must have $C^{*}\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(M))$, by Proposition 4.9. Preimages under $f$ of completely prime filters are completely prime. This means that preimages of filters in $C^{*}$ are in $\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$. By Theorem 6.1, there is a morphism $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})\rightarrow(M,C)$ which extends the frame map $f:L\rightarrow M$. ∎
6.4 Cofree Raney extension for frames with exact maps
Since for any Raney extension $(L,C)$ there is a subcolocale inclusion $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}$, there is a Raney surjection $(L,C)\to(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$. In light of Theorem 6.3 it is natural to wonder if $L\mapsto(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ is the object part of a right adjoint to $\pi_{1}:\mathbf{Raney}\to\mathbf{Frm}$. This is not the case. As shown in [ball19] – where the question is explored for structure $\mathsf{S}_{\mathfrak{c}}(L)$, isomorphic to $\mathsf{Filt}_{\mathcal{E}}(L)$ – not all frames maps $f:L\to M$ can be extended to their coframes of exact filters. For a frame morphism $f:L\rightarrow M$, we will say that it is exact if whenever the meet of a family $\{x_{i}:i\in I\}\subseteq L$ is exact, so is the meet of $\{f(x_{i}):i\in I\}$, and furthermore $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})$.
Proposition 6.6.
A morphism $f:L\rightarrow M$ is exact if and only if preimages of exact filters are exact. This holds if and only if the morphism can be extended to a Raney morphism
Proof.
Suppose that $f:L\rightarrow M$ is an exact frame map, and that $G\subseteq M$ is an exact filter. Suppose that $\bigwedge_{i}x_{i}\in L$ is an exact meet such that $f(x_{i})\in G$. By exactness of this map, the meet $\bigwedge_{i}f(x_{i})$ is exact and so $\bigwedge_{i}f(x_{i})\in G$. Again, by exactness of $f$, $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})$. Indeed, then, $\bigwedge_{i}x_{i}\in f^{-1}(G)$. Conversely, suppose that there is a frame map $f:L\rightarrow M$ that it is not exact. This means that either there is an exact meet $\bigwedge_{i}x_{i}\in L$ such that it is not preserved by $f$, or there is an exact meet $\bigwedge_{i}x_{i}\in L$ such that $\bigwedge_{i}f(x_{i})$ is not exact. We consider these two cases in turn. In the first case, we consider the principal filter ${\uparrow}\bigwedge_{i}f(x_{i})$. This is exact, as it is closed under all meets. We notice that by our hypothesis $f(\bigwedge_{i}x_{i})$ is not an element of this filter. Let us call $F$ the preimage of this filter. Both $x_{i}\in F$ and $\bigwedge_{i}x_{i}\notin F$, and so $F$ is not exact. In the second case, consider an exact meet $\bigwedge_{i}x_{i}\in L$ such that $\bigwedge_{i}f(x_{i})$ is not exact. In particular, let $y\in M$ be such that $\bigwedge_{i}(f(x_{i})\vee y)\nleq(\bigwedge_{i}f(x_{i}))\vee y$. We now consider the exact filter
That this is an exact filter follows from the characterization of Lemma 2.11. Let $F$ be the preimage of this filter. For each $i\in I$, $x_{i}\in F$. We claim that $\bigwedge_{i}x_{i}\notin F$. This follows from the fact that by our hypothesis $\bigwedge_{i}(f(x_{i})\vee y)\nleq(\bigwedge_{i}f(x_{i}))\vee y$ and $f(\bigwedge_{i}x_{i})\leq\bigwedge_{i}f(x_{i})$. The rest of the claim follows by Theorem 6.1. ∎
Let us call $\mathbf{Frm}_{\mathcal{E}}$ the category of frames with exact maps. Let us also call $\mathbf{Raney}_{\mathcal{E}}$ the category of Raney extensions with morphisms $f$ such that $\pi_{1}(f)$ is exact. The assignment $L\mapsto(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$, then, determines a functor $\mathsf{Filt}_{\mathcal{E}}:\mathbf{Frm}_{\mathcal{E}}\to\mathbf{Raney}_{\mathcal{E}}$.
Theorem 6.7.
There is an adjunction $\pi_{1}:\mathbf{Raney}_{\mathcal{E}}\leftrightarrows\mathbf{Frm}_{\mathcal{E}}:\mathsf{Filt}_{\mathcal{E}}$ with $\pi_{1}\dashv\mathsf{Filt}_{\mathcal{E}}$.
Proof.
Suppose that $f:L\rightarrow M$ is an exact frame map, and that $(L,C)$ is a Raney extension. By Proposition 6.6, as $f$ is exact, preimages of filters in $\mathsf{Filt}_{\mathcal{E}}(M)$ are in $\mathsf{Filt}_{\mathcal{E}}(L)$. Furthermore, $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}$ by Proposition 3.8. Then, preimages of exact filters of $M$ are in $C^{*}$. By Theorem 6.1, then, there is a map of Raney extensions $(L,C)\rightarrow(M,\mathsf{Filt}_{\mathcal{E}}(M)^{op})$ extending $f$, as desired. ∎
Recall that we have defined the Raney extensions of the form $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ as the $T_{D}$ ones. Let $\mathbf{Raney_{D}}$ be the full subcategory of $\mathbf{Raney}$ given by the $T_{D}$ Raney extensions.
Proposition 6.8.
The forgetful functor $\pi_{1}:\mathbf{Raney}\to\mathbf{Frm}$ restricts to an isomorphism $\mathbf{Frm}_{\mathcal{E}}\cong\mathbf{Raney_{D}}$.
Proof.
For a map $f:(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})\to(M,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ of $T_{D}$ Raney extensions, by Theorem 6.1 the restriction $f|_{L}:L\to M$ is a map in $\mathbf{Frm}_{\mathcal{E}}$. Thus, the restriction and co-restriction of $\pi_{1}$ is well-defined. The inverse functor maps each frame $L$ to the Raney extension $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$, and this assignment is functorial by Proposition 6.6. ∎
In contrast with the frame setting, the $T_{D}$ objects form a full subcategory of our pointfree category. We will explore the consequences of this in relation to the $T_{D}$ duality in Subsection 7.2
6.5 Sober coreflection of a Raney extension
Theorem 6.1, more broadly, can also be used to construct coreflections and reflections. For a Raney extension $(L,C)$ we call a map $\sigma:S(L,C)\rightarrow(L,C)$ of the category $\mathbf{Raney}$ a sobrification if $S(L,C)$ is sober, and if whenever $f:(M,D)\rightarrow(L,C)$ is a morphism from a sober Raney extension, there a commuting diagram
${S(L,C)}$${(L,C).}$${(M,D)}$$\scriptstyle{\sigma}$$\scriptstyle{f_{\sigma}}$$\scriptstyle{f}$
Theorem 6.9.
For a Raney extension $(L,C)$, the map
$\sigma:(L,\mathcal{I}(C^{*}\cup\mathsf{Filt}_{\mathcal{CP}}(L))^{op})\rightarrow(L,C)$
$F\mapsto\bigwedge F$
is its sobrification.
Proof.
Observe that, as $C^{*}\subseteq\mathsf{Filt}(L)$ is a sublocale and by Lemmas 4.17 and 4.18, $\mathcal{I}(C^{*}\cup\mathsf{Filt}_{\mathcal{CP}}(L))$ is a sublocale. As $(L,\mathcal{I}(C^{*}\cup\mathsf{Filt}_{\mathcal{CP}}(L))^{op})$ contains all completely prime filters of $L$, indeed, by Proposition 3.2 it is $\mathcal{CP}$-compact. Since $C^{*}\subseteq\mathcal{I}(C^{*}\cup\mathsf{Filt}_{\mathcal{CP}}(L))$, by Theorem 6.1 it means that the identity on $L$ extends to a surjective map of Raney extensions
$\sigma:(L,\mathcal{I}(C^{*}\cup\mathsf{Filt}_{\mathcal{CP}}(L))^{op})\to(L,C)$
$F\mapsto\bigwedge F.$
Let us show that this map has the required universal property. Suppose that $f:(M,D)\rightarrow(L,C)$ is a Raney map from a sober Raney extension. Consider the frame map $f|_{M}:M\rightarrow L$. By Theorem 6.1, to show that the map lifts it suffices to show that the preimage of each filter in $\mathsf{Filt}_{\mathcal{CP}}(L)$ as well as each filter in $C^{*}$ is in $D^{*}$. For filters in $C^{*}$, this holds because there is a map $f:(M,D)\rightarrow(L,C)$. For a completely prime filter $P\subseteq L$, recall that $f^{-1}(P)\in\mathsf{Filt}_{\mathcal{CP}}(M)$, and by definition of sobriety and Proposition 3.2, also $\mathsf{Filt}_{\mathcal{CP}}(M)\subseteq D^{*}$. ∎
6.6 TD reflection for Raney extensions and exact maps
We have seen that in $\mathbf{Raney}$, in contrast to $\mathbf{Frm}$, one can speak of $T_{D}$ objects that are not necessarily spatial. This enables us to define a notion of $T_{D}$ reflection of an object which does not spatialize said object. Once again, the morphisms have to be restricted. For a Raney extension $(L,C)$, we call a $T_{D}$ reflection a map $\delta:(L,C)\to D(L,C)$ such that $D(L,C)$ is $T_{D}$, and such that whenever $f:(L,C)\to(M,D)$ is a map to a $T_{D}$ Raney extension, these is a commuting diagram as follows.
${(L,C)}$${D(L,C)}$${(M,D).}$$\scriptstyle{\delta}$$\scriptstyle{f}$$\scriptstyle{f_{\delta}}$
Proposition 6.10.
In the category $\mathbf{Raney}_{\mathcal{E}}$, every Raney extension admits a $T_{D}$ reflection.
Proof.
We claim that the required map for $(L,C)$ is the surjection $\delta:(L,C)\to(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ following from Proposition 3.8. This is a map in $\mathbf{Raney}_{\mathcal{E}}$, as it restricts to the identity on $L$. Now, suppose that there is a $T_{D}$ Raney extension $(M,D)$ such that there is a morphism $f:(L,C)\to(M,D)$ in $\mathbf{Raney}_{\mathcal{E}}$. Because $(M,D)$ is $T_{D}$, by definition $D^{*}=\mathsf{Filt}_{\mathcal{E}}(L)^{op}$. By assumption on $f$, then, the preimage map relative to $f|_{L}$ maps filters in $D^{*}$ to exact filters of $L$. Hence, by Theorem 6.1, there is a map $f_{\delta}:(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})\to(M,D)$ as required. Finally, this map is in $\mathbf{Raney}_{\mathcal{E}}$ as it extends $f|_{L}$. ∎
6.7 Canonical extension as free algebraic Raney extension on a pre-spatial frame
We now view the canonical extension of a frame from [jakl20] as a Raney extension, and characterize it as a free construction. For a pre-spatial frame $L$, we will call its canonical extension the Raney extension $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})$. For this pair to be a Raney extension, we do need pre-spatiality, by Proposition 2.14.
@@ 6.7 Canonical extension as free algebraic Raney extension on a pre-spatial frame
Any frame morphism $f:L\rightarrow M$ between pre-spatial frames extends to a Raney morphism
$f_{\mathcal{SO}}:(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})\rightarrow(M,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(M))^{op}).$
@@
For a Raney extension $(L,C)$, we say that an element $c\in C$ is compact if, for every directed collection $D\subseteq L$, $c\leq\bigvee D$ implies that $c\leq d$ for some $d\in D$. We say that a Raney extension $(L,C)$ is algebraic if every element of $c$ is the join of compact elements.
@@
Notice that an element $x\in C$ is compact if and only if the filter ${\uparrow}^{L}x$ is Scott-open. Consider the isomorphism ${\uparrow}^{L}:C\cong C^{*}$. The Raney extension $(L,C)$ is algebraic if and only if in $C^{*}$ every element is a join of Scott-open filters of the form ${\uparrow}^{L}x$ for some $x\in L$. The inclusion $C^{*}\subseteq\mathsf{Filt}(L)^{op}$ is a subcolocale inclusion, and subcolocale inclusions preserves all joins, and joins in $\mathsf{Filt}(L)^{op}$ are intersections. Therefore, algebraicity of $(L,C)$ is equivalent to every filter in $C^{*}$ being an intersection of Scott-open filters in $C^{*}$. ∎
@@
For a pre-spatial frame $L$, its canonical extension is the free algebraic Raney extension over it.
@@
Suppose that $L$ is a pre-spatial frame, and that $(M,C)$ is an algebraic Raney extension. Suppose that there is a frame map $f:L\rightarrow M$. Consider the canonical extension $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})$. As $(M,C)$ is algebraic, $C^{*}\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(M))^{op}$, by Lemma 6.12. By Lemma 6.11, preimages of Scott-open filters are Scott-open. Then, preimages of filters in $C^{*}$ are in $\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$. By Theorem 6.1 this means that there is a map of Raney extensions $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})\rightarrow(M,C)$ extending the frame map $f:L\rightarrow M$. ∎
7 Special topics
7.1 Sobriety and strict sobriety
We work towards characterizing sobriety and strict sobriety of spaces in terms of Raney extensions.